A steel pipe of outside diameter 20 mm and thickness 3 mm is
deflected by 3 mm when used as a beam supported at its ends, 1
m apart, and subjected to a central load of 170 N. Find the
buckling load when the pipe is used as a column with hinged
ends. What is the maximum lateral deflection of this column
before the material attains the yield stress of 250N/mm2?
1
Expert's answer
2020-12-07T03:48:16-0500

a. Steel pipe is used as a simply supported beam.

The deflection formula for simply supported beam is "\sigma = \frac{PL^3}{48EI}"


"I = \frac{\pi}{64}*(D^4-d^4)= \frac{\pi}{64}*(20^4-14^4)=5968.2406 mm^4"


"Therefore 3 = \frac{170*1000^3}{48*E5968.2406}"


"E=197806.293 MPa = 197.806 MPa"


b. If the steel pipe is used as a column with both ends hinged


The buckling load for both ends, "P_b=\frac{\pi^2EI}{Le^2}"


"P_b=\frac{\pi^2*197806.329*5968.2406}{1000^2} = 11651.62 N"